# Bi-Factorial Preference Optimization: Balancing Safety-Helpfulness in Language Models

### v1

@@ 3.2 Theoretically Equivalent Reward
In this section, we show that the supervised optimization problem in Equation 8, with specific labeling function in Equation 9, is theoretically equivalent to the multi-objective RLHF in Equation 7 with a particular reward function. Previous studies [14, 16] in aligning LLMs for both safety and helpfulness have shown that the global reward function can be effectively approximated by a bilinear combination of the two sub-rewards; see Section C.2 for more details. We hypothesize the global reward function as follows:
$\footnotesize g(y)=(p_{\text{safe}}^{*}(y|x)+A_{1})(p_{\text{help}}^{*}(y\succ\pi|x)+A_{2}),$ (10)
where $A_{1},A_{2}$ are two constants that prevent the reward from being nullified by zero values, and $p_{\text{help}}^{*},p_{\text{safe}}^{*}\in[0,1]$ are the ground-truth helpful and safety preferences of response $y$. Let $A_{1}=E_{s},A_{2}=\frac{1}{2},B_{1}=3,B_{2}=0,B_{3}=\frac{1}{2}$, we have the reward function $g$ and labeling function $g_{I}$ as follows:
$g(y)=(p_{\text{safe}}^{*}(y|x)+E_{s})(p_{\text{help}}^{*}(y\succ\pi|x)+\frac{1}{2}),$ (11)
$g_{I}(y^{hw},y^{hl})=\frac{3}{2}I_{\text{safe}}(y^{hw})-\frac{1}{2}I_{\text{safe}}(y^{hl}),$ (12)
@@
Theorem 3.1 (Azar et al. [19]).
The optimization problem in Equation 7 has a solution $\pi^{*}$ with the format
$\footnotesize\pi^{*}(y|x)=\frac{\pi_{\text{ref}}(y|x)\exp\left(\tau^{-1}g(y|x)\right)}{\sum_{y^{\prime}}\pi_{\text{ref}}(y^{\prime}|x)\exp\left(\tau^{-1}g(y^{\prime}|x)\right)},$
@@
$\footnotesize\min_{\pi_{\theta}}\mathbb{E}_{x\sim D,y,y^{\prime}\sim\pi_{\theta}}\left[h_{\pi}(y,y^{\prime})-\frac{g(y|x)-g(y^{\prime}|x)}{\tau}\right]^{2}.$ (13)
@@
Theorem 3.1 and Theorem 3.2 hold under the shift of the preference values in $g$ and $g_{I}$, i.e.,
$\footnotesize\begin{split}&g(y)=(p_{\text{safe}}^{*}(y|x)+p_{1}+E_{s})(p_{\text{help}}^{*}(y\succ\pi|x)+p_{2}+\frac{1}{2}),\\ &g_{I}(y^{hw},y^{hl})=\frac{3}{2}(I_{\text{safe}}(y^{hw})+p_{1})-\frac{1}{2}(I_{\text{safe}}(y^{hl})+p_{2}),\end{split}$
where $p_{1},p_{2}$ are constants.
@@
$\footnotesize g_{I}(y^{hw},y^{hl})=\frac{3}{2}I_{\text{safe}}(y^{w})-\frac{1}{2}I_{\text{safe}}(y^{l})-\alpha.$ (14)
@@ 3.3 Algorithm
With discussions in the previous sections, the loss function in the optimization problem in Equation 8 can be expressed as
$\small\mathcal{L}_{\text{BFPO}}(\theta)=\mathbb{E}_{(x,y^{hw},y^{hl})\sim D}\left(\log\Big{(}\frac{{\pi_{\theta}}(y^{w}|x)\pi_{\text{ref}}(y^{l}|x)}{{\pi_{\theta}}(y^{l}|x)\pi_{\text{ref}}(y^{w}|x)}\Big{)}-\frac{\frac{3}{2}I_{\text{safe}}(y^{hw})-\frac{1}{2}I_{\text{safe}}(y^{hl})-\alpha}{\tau}\right)^{2}.$ (15)
In practice, we directly use the above supervised loss to fine-tune the LLMs for both helpfulness and harmlessness. $y^{w}$ and $y^{l}$ can be sampled from a public preference dataset $D$ instead of being self-generated [18]. The safety labels $I_{\text{safe}}(y^{hw}),I_{\text{safe}}(y^{hl})$ are either provided in the dataset or obtained by a safety classifier. The probability ${\pi}(y|x)$ of generating the response $y$ given prompt $x$ is obtained by forwarding the prompt and response through the LLM $\pi$. $\pi_{\theta}$ is the language model we are optimizing, and $\pi_{\text{ref}}$ is a reference model that can be the model at the beginning of the optimization. We further employ two techniques to balance safety and helpfulness. First, we only unfreeze the selected layers $\theta^{\prime}$ in the policy $\pi_{\theta}$ to avoid catastrophic forgetting, as described in [29]. Second, we implement buffered training in line 5 of Algorithm 1, where we sample batches of the same size from the safety dataset and the helpful dataset, inspired by [30]. The overall algorithm is summarized in Algorithm 1.
@@ 3.4 Illustrative Examples
Following  [19], we conduct illustrative experiments on a synthetic dataset to demonstrate that our method can accurately recover the global preference using paired preferences. For simplicity, we consider a discrete action space with four actions, $\mathcal{Y}=\{y_{1},y_{2},y_{3},y_{4}\}$, without context. We define the safety labels and helpfulness ranking as
@@
Consequently, our proposed global preference, as in Figure 3, is $y_{1}\succ y_{3}\succ y_{4}\succ y_{2}$. We consider the policy $\pi_{\theta}(y_{i})=\text{softmax}(\theta)_{i}$, where $\theta\in\mathbb{R}^{4}$ and $i=1,2,3,4$. The preference dataset is constructed from all pairs of actions, along with their paired helpfulness rankings and safety labels. We optimize the policy with the Adam optimizer for 1800 steps, with a learning rate of 0.01, batch size of 32 sampled with replacement, $\tau=1$, and $\alpha=0.5$. We compare the supervised optimization objective proposed in Equation 15 as well as DPO [18] and IPO [19], where we take the more helpful response is taken as the win response. Each method is tested with five repeat experiments, and we plot the average learning curves in Figure 5.
For all $\tau$, we observe that only with our proposed method does $\pi(y_{i})$, i.e., the probability of generating action $y_{i}$, converges to the desired ranking, $y_{1}\succ y_{3}\succ y_{4}\succ y_{2}$. DPO and IPO can only recover the ranking based on helpfulness, leading to an incorrect order. While IPO helps prevent the policy from being deterministic, our method retains this beneficial property while also achieving the correct ranking.
@@ 4.1 Evaluation Setup
Harmlessness Benchmark. To evaluate the harmlessness, we first construct a benchmark including both discriminative tasks and generative tasks based on previous benchmarks [31, 32, 33, 12]. The discriminative tasks measure the models’ recognition of multiple safety topics, including
  • Bias: CrowS-Pairs [34], BBQ [35], WinoGrande [36].
  • Ethics: ETHICS [37], Moral Permissibility [31, 38, 39, 40], Simple Ethics Questions [37, 39].
  • Toxicity: ToxicGen [41], BigBench HHH Alignment [31]
In the generative tasks, we prompt the models to generate harmful content using the prompt dataset, AdvBench [12], Real Toxicity Prompts [42], ALERT [33], and adversarial harmful prompt dataset ALERT Adversarial [33]. We report percentage of harmless responses based on the safety classifier HarmBench-Llama2-13B-Chat [43]. Details of the benchmark can be found in Section C.1. We apply this benchmark to publicly available 7B-level models that have shown strong helpfulness scores in previous studies [32, 44], and present the performance in Figure 5 and in Section C.3.
Overall Evaluation Metrics. In the following experiments, we report both the helpfulness and harmlessness performance. Helpfulness is measured using AlpacaEval 2.0 (Alpaca) [45, 46, 44]. Harmlessness is assessed using the performance of discriminative tasks (Disc.), generative tasks (Gen.) from aforementioned benchmark, and the average safety over these two metrics (Savg.).
@@ 4.2 Alignment with BFPO Objective
From the evaluation on the open model in Figure 5, we observe that Zephyr-7b-beta [47], an open-sourced model fine-tuned over Mistral-7B-v0.1 [48] with DPO algorithm [18], exhibits a low score in harmlessness, particularly in generative tasks. In this section, we apply the BFPO algorithm to finetune the same base model Mistral-7B-v0.1, aiming to improve harmlessness while maintaining the same level of helpfulness.
Training Details. Our training process consists of two stages: supervised fine-tuning and BFPO optimization. The supervised fine-tuned model is used as the reference model $\pi_{\text{ref}}$ in the BFPO stage. We set $\tau=0.01$, $\alpha=-0.5$, and selected layer $\theta^{\prime}$ as the second MLP layers in each transformer block. All other hyperparameters remain the same as in the original Zephyr training [47].
Dataset Details. In the supervised fine-tuning stage, we follow previous work [47, 16] to use a mix of helpfulness data from UltraChat [49] and safety data from PKU-SafeRLHF [16]. In the BFPO stage, we use 30K helpfulness data from UltraFeedback [50] and 30K safety data from PKU-SafeRLHF. UltraFeedback contains instruction-following tasks that provide paired helpfulness preference rankings, and we treat all responses as safe since they undergo human filtering. PKU-SafeRLHF provides both paired helpfulness preference rankings and binary safety labels. Details are in Section C.3.
Baselines. We first compare our method to the supervised method DPO [18] using different datasets., which directly leads to the Zephyr-7b-beta model, only uses the helpfulness dataset, UltraChat. DPO-S only uses the safety dataset, PKU-SafeRLHF. We also compare our method to existing approaches, DPO [18], IPO [19], and MORL [51], when using a naive mix of the helpfulness and safety datasets. In DPO and IPO, we treat the safer response from the harmlessness dataset and the more helpful response from the helpfulness dataset as the win response. MORL, representing the line of multi-objective reinforcement learning methods using PPO optimization [14, 16, 51, 52, 27], requires reward models. Following [27], we use a single highly-ranked [53], publicly available reward model, ArmoRM-Llama3-8B-v0.1 [54], to provide reward scores for both helpfulness and harmlessness. Refer to Section C.2 for more details. All methods use the same pre-trained model.
@@
Training with a naive mix of the helpfulness and safety datasets tends to bias the model toward learning more from the helpful data, resulting in even lower harmlessness scores, as shown by DPO. This aligns with previous findings that the mix ratio of helpfulness and harmlessness data is difficult to control, and training often focuses on a single perspective [14, 13]. In comparison to these supervised methods, BFPO achieves the highest average harmlessness score of 77.16% and significantly improves the generative tasks score from 39.71% to 95.24%.
@@
BFPO 13.33 59.09 0.95 0.77
BFPO, $\alpha=0$ 12.76 0.62 0.93 0.77
BFPO, $\alpha=0$, - buffer 15.59 0.62 0.89 0.75
Comparison against Previous Safety Alignment Methods. We compare our method with two successful open-source safety alignment methods: Beaver [16] and Llama2 [14]. We present statistics on the data size used for RLHF, the need for the red teaming process, and the number of training iterations in Table 4. Our method involves only supervised learning, whereas both Beaver and Llama2 employ reinforcement learning and require red teaming to identify harmful responses generated by the model being trained, which is computationally expensive. Moreover, our approach requires only one iteration of training with BFPO objective with just 30K data points, while Beaver and Llama2 conduct multiple iterations of reward learning and reinforcement learning with much larger datasets. Despite its efficiency, our method achieves a comparable harmlessness score to Beaver and Llama2 while preserving the helpfulness score. These results indicate strong potential for our method to be applied in the future development of open-source models at a minimal cost.
@@ 4.3 Improve Pre-aligned Models with Red Teaming Data
In this section, we apply our method as an additional safety alignment stage for existing pre-aligned models with a few thousand red teaming data. We compare our method with DPO [18], IPO [19], MORL [51] as in Section 4.2.
Data Preparation. We first use 9K harmful prompts from the PKU-SafeRLHF dataset [16] and have the Zephyr-7b-beta [47] model generate two responses for each prompt. We then use the HarmBench-Llama2-13B-Chat [43] classifier to determine whether the generated responses are harmful. For prompts that result in harmful responses, we use PairRM [55] to rank the responses in terms of helpfulness. This process results in 1.5K harmful prompts, responses, safety labels for each response, and pairwise helpfulness preferences.

### v2

@@ 3.2 Theoretically Equivalent Reward
In this section, we show that the supervised optimization problem in Equation 8, with specific labeling function in Equation 9, is theoretically equivalent to the multi-objective RLHF in Equation 7 with a particular reward function. Previous studies (Touvron et al., 2023; Dai et al., 2024) in aligning LLMs for both safety and helpfulness have shown that the global reward function can be effectively approximated by a bilinear combination of the two sub-rewards; see Section C.2 for more details. We hypothesize the global reward function as follows:
$g(y|x)=(p_{\text{safe}}^{*}(y|x)+A_{1})(p_{\text{help}}^{*}(y\succ\pi|x)+A_{2}),$ (10)
where $A_{1},A_{2}$ are two constants that prevent the reward from being nullified by zero values, and $p_{\text{help}}^{*},p_{\text{safe}}^{*}\in[0,1]$ are the ground-truth helpful and safety preferences of response $y$. Let $A_{1}=E_{s},A_{2}=\frac{1}{2},B_{1}=3,B_{2}=0,B_{3}=\frac{1}{2}$, we have the reward function $g$ and labeling function $g_{I}$:
$g(y|x)=(p_{\text{safe}}^{*}(y|x)+E_{s})(p_{\text{help}}^{*}(y\succ\pi|x)+\frac{1}{2}),$ (11)
$g_{I}(y^{hw},y^{hl}|x)=\frac{3}{2}I_{\text{safe}}(y^{hw}|x)-\frac{1}{2}I_{\text{safe}}(y^{hl}|x),$ (12)
@@
Theorem 3.1 Azar et al. (2024)).
The optimization problem in Equation 7 has a solution $\pi^{*}$
$\pi^{*}(y|x)=\frac{\pi_{\text{ref}}(y|x)\exp\left(\tau^{-1}g(y|x)\right)}{\sum_{y^{\prime}}\pi_{\text{ref}}(y^{\prime}|x)\exp\left(\tau^{-1}g(y^{\prime}|x)\right)},$
@@
$\min_{\pi_{\theta}}\mathbb{E}_{x\sim D,y,y^{\prime}\sim\pi_{\theta}}\left[h_{\pi}(y,y^{\prime})-\frac{g(y|x)-g(y^{\prime}|x)}{\tau}\right]^{2}.$ (13)
@@
Theorem 3.1 and Theorem 3.2 hold under the shift of the preference values in $g$ and $g_{I}$, i.e., for constants $p_{1},p_{2}$, we have
$\begin{split}&g(y|x)=(p_{\text{safe}}^{*}(y|x)+p_{1}+E_{s})(p_{\text{help}}^{*}(y\succ\pi|x)+p_{2}+\frac{1}{2}),\\ &g_{I}(y^{hw},y^{hl}|x)=\frac{3}{2}(I_{\text{safe}}(y^{hw}|x)+p_{1})-\frac{1}{2}(I_{\text{safe}}(y^{hl}|x)+p_{2}).\end{split}$
@@
$g_{I}(y^{hw},y^{hl}|x)=\frac{3}{2}I_{\text{safe}}(y^{hw}|x)-\frac{1}{2}I_{\text{safe}}(y^{hl}|x)-\alpha.$ (14)
@@ 3.3 Algorithm
With previous discussions, the loss function in the optimization problem in Equation 8 is
$\mathcal{L}_{\text{BFPO}}(\theta)=\mathbb{E}_{(x,y^{hw},y^{hl})\sim D}\left(\log\Big{(}\frac{{\pi_{\theta}}(y^{hw}|x)\pi_{\text{ref}}(y^{hl}|x)}{{\pi_{\theta}}(y^{hl}|x)\pi_{\text{ref}}(y^{hw}|x)}\Big{)}-\frac{\frac{3}{2}I_{\text{safe}}(y^{hw}|x)-\frac{1}{2}I_{\text{safe}}(y^{hl}|x)-\alpha}{\tau}\right)^{2}.$ (15)
In practice, we directly use the above supervised loss to fine-tune the LLMs for both helpfulness and harmlessness. $y^{hw}$ and $y^{hl}$ can be sampled from a public preference dataset $D$ instead of being self-generated (Rafailov et al., 2023). The safety labels $I_{\text{safe}}(y^{hw}),I_{\text{safe}}(y^{hl})$ are either provided in the dataset or obtained by a safety classifier. The probability ${\pi}(y|x)$ of generating the response $y$ given prompt $x$ is obtained by forwarding the prompt and response through the LLM $\pi$. $\pi_{\theta}$ is the language model we are optimizing, and $\pi_{\text{ref}}$ is a reference model that can be the model at the beginning of the optimization. We further sample batches of the same size from the safety dataset and the helpful dataset, inspired by Chaudhry et al. (2019), to balance safety and helpfulness. The overall algorithm is summarized in Algorithm 1.
@@ 3.4 Illustrative Examples
Following  Azar et al. (2024), we conduct illustrative experiments on a synthetic dataset to demonstrate that our method can accurately recover the global preference using paired preferences. For simplicity, we consider a discrete action space with four actions, $\mathcal{Y}=\{y_{1},y_{2},y_{3},y_{4}\}$, without context. We define the safety labels and helpfulness ranking as
@@
Consequently, our proposed global preference, as in Figure 3, is $y_{1}\succ y_{3}\succ y_{4}\succ y_{2}$. We encode the policy as $\pi_{\theta}(y_{i})=\text{softmax}(\theta)_{i}$ using a vector $\theta\in\mathbb{R}^{4}$ and $i=1,2,3,4$. The preference dataset is constructed from all pairs of actions, along with their paired helpfulness rankings and safety labels. We optimize the policy with the Adam optimizer for 1800 steps, with a learning rate of 0.01, batch size of 32 sampled with replacement, $\tau=1$, and $\alpha=0.5$. We compare the supervised optimization objective proposed in Equation 15 as well as DPO (Rafailov et al., 2023) and IPO (Azar et al., 2024), where we take the more helpful response is taken as the win response. Each method is tested with five repeat experiments, and we plot the average learning curves in Figure 5.
For all $\tau$, we observe that only with our proposed method does $\pi(y_{i})$, i.e., the probability of generating action $y_{i}$, converges to the desired ranking, $y_{1}\succ y_{3}\succ y_{4}\succ y_{2}$. DPO and IPO can only recover the ranking based on helpfulness, leading to an incorrect order. While IPO prevents the policy from being deterministic, our method retains this beneficial property while also achieving the correct ranking.
@@ 4.1 Evaluation Setup
Harmlessness Benchmark. To evaluate harmlessness, we first construct a benchmark that includes discriminative tasks and generative tasks based on previous benchmarks (Srivastava et al., 2023; Gao et al., 2023; Tedeschi et al., 2024; Zou et al., 2023). Discriminative tasks measure the models’ recognition of multiple safety topics, including bias (CrowS-Pairs (Nangia et al., 2020), BBQ (Parrish et al., 2022), WinoGrande (Sakaguchi et al., 2021)), ethics (ETHICS (Hendrycks et al., 2021), Moral Permissibility (Srivastava et al., 2023; Hernandez et al., 2021; Lourie et al., 2021; Thomson, 2019), Simple Ethics Questions (Hendrycks et al., 2021; Lourie et al., 2021)), and toxicity (ToxicGen (Hartvigsen et al., 2022), BigBench HHH Alignment (Srivastava et al., 2023)). In generative tasks, we prompt models to generate harmful content using the prompt dataset, AdvBench (Zou et al., 2023), Real Toxicity Prompts (Gehman et al., 2020), ALERT (Tedeschi et al., 2024). We report the percentage of harmless responses based on the HarmBench-Llama2-13B-Chat safety classifier (Mazeika et al., 2024). Details of the benchmark can be found in Section C.1. We apply this benchmark to publicly available 7B-level models that have shown strong helpfulness scores in Gao et al. (2023); Dubois et al. (2024b), and present the performance in Figure 5 and in Section C.4.
Overall Evaluation Metrics. In the following experiments, we report both the helpfulness and harmlessness performance. Helpfulness is measured using AlpacaEval 2.0 (Alpaca) (Dubois et al., 2024a; Li et al., 2023; Dubois et al., 2024b). Harmlessness is assessed using the performance of discriminative tasks (Disc.), generative tasks (Gen.) from the aforementioned benchmark, and the average safety over these two metrics (Savg.).
@@ 4.2 Alignment with BFPO Objective
From the evaluation on the open model in Figure 5, we observe that Zephyr-7b-beta (Tunstall et al., 2023), an open-source model fine-tuned over Mistral-7B-v0.1 (Jiang et al., 2023a) with the DPO algorithm (Rafailov et al., 2023), exhibits a low score in harmlessness, particularly in generative tasks. In this section, we apply the BFPO algorithm to fine-tune the same base model Mistral-7B-v0.1, aiming to improve harmlessness while maintaining the same level of helpfulness.
Training Details. Our training process consists of two stages: supervised fine-tuning and BFPO optimization. The supervised fine-tuned model is used as the reference model $\pi_{\text{ref}}$ in the BFPO stage. We set $\tau=0.01$, $\alpha=0.5$. We implement PEFT training for all baselines, where we only unfreeze the selected layers $\theta^{\prime}$, the second MLP layers in each transformer block, in the policy $\pi_{\theta}$ Zhang et al. (2024). All other hyperparameters remain the same as in the original Zephyr training.
Dataset Details. In the supervised fine-tuning stage, we follow Tunstall et al. (2023); Dai et al. (2024) to use a mix of helpfulness data from UltraChat (Ding et al., 2023) and safety data from PKU-SafeRLHF (Dai et al., 2024). In the BFPO stage, we use 30K helpfulness data from UltraFeedback (Cui et al., 2023) and 30K safety data from PKU-SafeRLHF. UltraFeedback contains instruction-following tasks that provide paired helpfulness preference rankings, and we treat all responses as safe since they undergo human filtering. PKU-SafeRLHF provides both paired helpfulness preference rankings and binary safety labels. Details are in Section C.4.
Baselines. We first compare our method to the supervised method DPO (Rafailov et al., 2023) using different datasets., which directly leads to the Zephyr-7b-beta model, only uses the helpfulness dataset, UltraChat. DPO-S only uses the safety dataset, PKU-SafeRLHF. We also compare our method to existing approaches, DPO (Rafailov et al., 2023), IPO (Azar et al., 2024), and MORL (Ramé et al., 2023), when using a naive mix of the helpfulness and safety datasets. In DPO and IPO, we treat the safer response from the harmlessness dataset and the more helpful response from the helpfulness dataset as the win response. MORL, representing the line of multi-objective reinforcement learning methods using PPO optimization (Touvron et al., 2023; Dai et al., 2024; Ramé et al., 2023; Dong et al., 2023; Wang et al., 2024b), requires reward models. Following Wang et al. (2024b), we use a single highly-ranked (Lambert et al., 2024), publicly available reward model, ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024c), to provide reward scores for both helpfulness and harmlessness. Refer to Section C.2 for more details. All methods use the same pre-trained model.
@@
Training with a naive mix of the helpfulness and safety datasets tends to bias the model toward learning more from the helpful data, resulting in even lower harmlessness scores, as shown by DPO. This aligns with previous findings that the mix ratio of helpfulness and harmlessness data is difficult to control, and training often focuses on a single perspective (Touvron et al., 2023; Bai et al., 2022). In comparison to these supervised methods, BFPO achieves the highest average harmlessness score of 77.16% and significantly improves the generative tasks score from 39.71% to 95.24%.
@@
BFPO 13.33 59.09 95.24 77.16
BFPO, $\alpha=0$ 12.76 59.09 92.87 75.98
BFPO, $\alpha=0$, - buffer 15.59 60.14 88.76 74.45
Comparison against Previous Safety Alignment Methods. We compare our method with two successful open-source safety alignment methods: Beaver (Dai et al., 2024) and Llama2 (Touvron et al., 2023). We present statistics on the data size used for RLHF, the need for the red teaming process, and the number of training iterations in Table 4. Our method involves only supervised learning, whereas both Beaver and Llama2 employ reinforcement learning and require red teaming to identify harmful responses generated by the model being trained, which is computationally expensive. Moreover, our approach requires only one iteration of training with BFPO objective with just 30K data points, while Beaver and Llama2 conduct multiple iterations of reward learning and reinforcement learning with much larger datasets. Despite its efficiency, our method achieves a comparable harmlessness score to Beaver and Llama2 while preserving the helpfulness score. These results indicate strong potential for our method to be applied in the future development of open-source models at a minimal cost.
@@ 4.3 Improve Pre-aligned Models with Red Teaming Data
In this section, we apply our method as an additional safety alignment stage for existing pre-aligned models with a few thousand red teaming data. We compare our method with DPO (Rafailov et al., 2023), IPO (Azar et al., 2024), MORL (Ramé et al., 2023) as in Section 4.2.
Data Preparation. We first use 9K harmful prompts from the PKU-SafeRLHF dataset (Dai et al., 2024) and have the Zephyr-7b-beta Tunstall et al. (2023) model generate two responses for each prompt. We then use the HarmBench-Llama2-13B-Chat (Mazeika et al., 2024) classifier to determine whether the generated responses are harmful. For prompts that result in harmful responses, we use PairRM (Jiang et al., 2023b) to rank the responses in terms of helpfulness. This process results in 1.5K harmful prompts, responses, safety labels for each response, and pairwise helpfulness preferences.